# Orbital Description of Landau Levels

### v2

@@ Abstract
We construct a minimal lattice model to provide an orbital description of lowest and first Landau levels. With the maximally localized Wannier functions with $s$, $p_{-}$, $p_{+}$ orbital characteristics, a three-orbital model is developed, where the lowest two Chern bands are flat with $\mathcal{C}=1$. This model can be viewed as consecutive band inversions between these Wannier states at $\Gamma$ and $K$ in momentum space, which adiabatically connects the atomic insulator limit to Landau level physics. Interestingly, many-body exact diagonalization and entanglement spectrum analysis suggest that the Abelian states can appear in the 1/3-filled lowest Chern band, while the signatures of the non-Abelian states are found in the half-filled first Chern band. This construction can be further extended to realize flat Chern bands resembling the higher Landau levels. Our results offer a new perspective to understand the lattice analogue of Landau levels, potentially enabling the realization of the fascinating topological phenomena at higher temperatures.
@@
Introduction.— The recent discovery of the Abelian fractional Chern insulators (FCI) at zero magnetic field, dubbed fractional quantum anomalous Hall state in moiré MoTe2 and pentalayer graphene [1, 2, 3, 4, 5] has generated intense interest in this new state of matter [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. The emergence of the fractional topological states is attributed to the existence of flat Chern bands [20, 21, 22, 23, 24, 25] with nearly ideal quantum geometry in moiré superlattice, which resembles the lowest Landau level (LLL). The exotic non-Abelian topological orders [26], such as Moore-Read states [27, 28, 29, 30] or Read-Rezayi state [31] are predicted to exist in the partially filled first Landau level (1LL). These states support non-Abelian quasiparticle excitations [28] and could be utilized as a platform for fault-tolerant quantum computation [32, 33]. This raises an interesting question [34, 35, 36, 37, 38, 39]: Is it possible and how to construct a flat Chern band on the lattice which resembles the 1LL physics with non-Abelian statistics? A positive answer to this question could offer opportunity to realize non-Abelian fractionalization at elevated temperatures without magnetic field.
Here we construct a minimal lattice model with maximally localized Wannier functions (MLWFs) with $s$, $p_{-}$, $p_{+}$ orbital characteristics. The lowest two bands are flat with Chern number equal to $1$, where they have band inversion at $\Gamma$, while the second lowest and highest bands invert at $K$. Interestingly, many-body exact diagonalization calculations are carried out with the Wannier functions, which suggests that 1/3-filled lowest Chern band (LCB) supports the Abelian states, while half-filled first Chern band (1CB) supports the non-Abelian states. The quantum geometry of the combined lowest two Chern bands is close to the ideal value of the combined lowest two LLs. Thus, the lowest two Chern bands constitute the lattice analogue of the LLL and 1LL. Our method can be further extended to realize flat Chern bands resembling the higher LLs, as well as cold atom systems.
@@
Intuition.— As a starting point, to construct the Chern bands in lattice model, we should know the characteristics of the Wannier states. Previously, Qi constructed the one-dimensional MLWF in Chern bands that have a one-to-one mapping to the LLL wave functions in quantum Hall [23] of the form $\psi_{0,K_{y}}(x,y)\sim e^{iK_{y}y}H_{0}(x^{\prime})e^{-x^{\prime 2}/2l^{2}_{B}}$, where $x^{\prime}\equiv x-K_{y}\ell_{B}^{2}$. $H_{0}\left(x-K_{y}\ell_{B}^{2}\right)$ is the Hermite polynomial for $n=0$ LLL, and $H_{0}(x^{\prime})e^{-x^{\prime 2}/2l^{2}_{B}}=e^{-x^{\prime 2}/2l^{2}_{B}}$ is related to MLWF of Chern band and is $s$ orbital like in one dimension. Intuitively, when the construction of more generic wave functions of Chern bands generalizes to 1LL, the MLWF of Chern bands maps to $H_{1}(x^{\prime})e^{-x^{\prime 2}/2l^{2}_{B}}=2x^{\prime}e^{-x^{\prime 2}/2l^{2}_{B}}$, which resembles $p$ orbital. Intuitively, at least the Wannier states with orbital characteristics $s$ and $p$ are needed to construct the Chern band that resembles 1LL physics.
@@
Model and adiabatic connection.— We begin with a minimal continuum model in which itinerant electrons couple to a layer pseudospin skyrmion lattice. This model could well describe the consecutive topologically nontrivial flat bands in twisted homobilayer MoTe2 around twist angle $2^{\circ}$ [40, 39, 41, 42, 43, 44, 45],
@@
Here $\bm{\sigma}$ is the pseudospin operator representing the degree of freedom of the layer, $\mathbf{S}(\mathbf{r})=\mathbf{S}(\mathbf{r}+\mathbf{a}_{1,2})$ represents the periodic moiré potential and is naturally coupled to the pseudospin of the layer with strength $\mathcal{J}$. When the twist angle is small, the moiré potential dominates $\mathcal{J}\gg\hbar^{2}/ma^{2}$ and enforces local electron layer polarization align with $\mathbf{S}(\mathbf{r})$, which in turn induces a pseudospin Berry phase. This can be seen explicitly by a position-dependent $SU(2)$ unitary transformation $\mathbf{U}(\mathbf{r})$ that rotates the pseudo-spin texture $\mathbf{S}(\mathbf{r})$ into $S(\mathbf{r})\hat{z}$ and introduces a gauge field $\mathbf{A}_{j}=(i\hbar/e)\mathbf{U}^{\dagger}\partial_{j}\mathbf{U}$. For large $\mathcal{J}$, we can further project the Hamiltonian onto the low energy manifold of local layer polarized electrons and obtain the effective Hamiltonian for the low energy band as
@@
From such observations, we see that $\mathcal{H}_{\text{eff}}$ actually links effective atomic orbital and LL physics in the low energy manifold. This is made clearly by calculating the band structure of $\mathcal{H}_{0}$ when we take $\mathbf{S}(\mathbf{r})=\mathbf{N}(\mathbf{r})/N^{\lambda}(\mathbf{r})$ with
@@
where $\hat{\mathbf{e}}_{j}=\left(i\alpha\sin\theta_{j},-i\alpha\cos\theta_{j},-1\right)/\sqrt{2}$, $\mathbf{q}_{j}=(4\pi/\sqrt{3}a)\left(\cos\theta_{j},\sin\theta_{j}\right)$, and the angles satisfy $\theta_{2}=\theta_{1}+2\pi/3$, $\theta_{3}=\theta_{1}+4\pi/3$, and $\theta_{j+3}=\theta_{j}+\pi$. This pseudospin texture can be viewed as a sum of three pseudospin spirals forming a triangular skyrmion lattice [49, 50, 51, 52], and the normalization is controlled by $\lambda$. When $\lambda=1$, $S(\mathbf{r})$ becomes spatially uniform, and the low energy bands of $\mathcal{H}_{0}$ are flat Chern bands with $\mathcal{C}=1$, as shown in Fig. 1(e). The lowest two Chern bands are dispersive LL, where their wave functions overlap with the corresponding flat LL exceed $99\%$ [53]. For simplicity, we label them as LLL and 1LL, respectively. Conversely, when $\lambda\rightarrow 0$, the minimum of $-JS(\mathbf{r})$ forms a moiré triangular lattice (see Fig. 1(g)), and three lowest bands in Fig. 1(a) are topologically trivial flat bands resembling the atomic insulator limit, which emerge from $s$, $p_{-}$, $p_{+}$ orbital hopping on a triangular lattice.
@@
Figure 1: Band structures and adiabatic connection from atomic insulator to LL. (a)-(e) The lowest few bands of $\mathcal{H}_{0}$ with different normalization parameter $\lambda$, where the overlap between Bloch state and Wannier function is labeled as different colors, and $\mathcal{C}$ denotes the Chern number of lowest three bands. The inset in (c) shows the Brillouin zone (BZ). (f) Band structure of the minimal three-orbital tight-binding model. (g) Real space distribution of $-N(\mathbf{r})$. The parameters are $J/(\hbar^{2}/ma^{2})=52\pi^{2}$, $\alpha=1$ and $N_{0}=0.28$.
@@
Fig. 1 display the evolution of the band structures from flat topologically trivial bands to flat Chern bands. As $\lambda$ changes from 0 to 1, the lowest three bands invert while their $s$, $p_{-}$, $p_{+}$ orbital characteristics remain. To see the band inversion clearly, we graphically depict the overlap of the Bloch states with trial Wannier functions of $s$, $p_{-}$, and $p_{+}$ orbitals, symbolized by blue, green, and red, respectively. The lowest three bands first invert with the upper band at $M$ and their Chern numbers become $\mathcal{C}=[0,0,3]$ as shown in Fig. 1(b), then the second and third lowest bands invert at $K$ and switch Chern numbers to $\mathcal{C}=[0,2,1]$ as in Fig. 1(c), finally the lowest and second lowest bands invert at $\Gamma$ and Chern numbers becomes $\mathcal{C}=[1,1,1]$ as in Fig. 1(d,e). The lowest three Chern bands are isolated in Fig. 1(e). The orbital characteristics of the LLL and 1LL here are consistent with the mapping from the hybrid Wannier functions of Chern band to LL wave functions in quantum Hall [23].
@@
Wannier projection.— Now we demonstrate that $s$, $p_{-}$, $p_{+}$ orbitals constitute a complete subspace to describe LLL and 1LL. By constructing the MLWFs [54, 55, 56, 57] from the continuum model, a three-orbital model is developed to describe the lowest two Chern bands in Fig. 1(e), with the Wannier orbitals forming a triangular lattice. The minimal model includes hopping up to the fifteenth nearest neighbors, where its explicit form is in Supplemental Materials [53]. The band structure in Fig. 1(e) does not possess a local gap below which the total Chern number is zero. A set of frozen states is chosen to preserve the topology of the focused Chern bands and the band disentanglement [58] process is then performed to avoid the Wannier obstruction [59]. We choose the lowest two Chern bands as frozen states, and get the minimal tight-binding model where only MLWFs with $s$, $p_{-}$, $p_{+}$ orbitals characteristics are included. The corresponding band structure is calculated in Fig. 1(f). The MLWFs of $s$, $p_{-}$, $p_{+}$ orbitals are shown in Fig. 2. The lowest two bands projected from the LLL and 1LL are well reconstructed in Fig. 1(f), which are labeled the LCB and 1CB, respectively.
@@
Figure 2: The MLWFs of the lowest few bands in Fig. 1(e) with $s$, $p_{-}$, $p_{+}$ orbital characteristics, and the while hexagon labels Wigner-Seitz cell.
@@
[…]
@@
where $\mathcal{B}(\mathbf{k})\equiv-2\mathrm{Im}(\eta^{xy})$ is the Berry curvature, $g(\mathbf{k})\equiv\mathrm{Re}(\eta^{\mu\nu})$ is the Fubini-Study metric, and $\eta^{\mu\nu}(\mathbf{k})\equiv\langle\partial^{\mu}u_{\mathbf{k}}|\left(1-|u_{\mathbf{k}}\rangle\langle u_{\mathbf{k}}|\right)|\partial^{\nu}u_{\mathbf{k}}\rangle$ is the quantum geometric tensor, $\mathcal{C}\equiv(1/2\pi)\int d^{2}\mathbf{k}\mathcal{B}(\mathbf{k})$, $\Omega_{\text{BZ}}$ is area of BZ. We plot the distribution of $\mathcal{B}(\mathbf{k})$ and $\text{Tr}[g(\mathbf{k})]$ of the BZ in Fig. 3. For LLL and 1LL, $\mathcal{B}(\mathbf{k})$ remain positive throughout the whole BZ and the distribution is quite homogenous with relatively small fluctuation $\delta\mathcal{B}$, and $\mathbb{T}$ is almost ideal ($\mathbb{T}=2n+1$ for $n$th flat LL). However, in the tight-binding model, $\mathcal{B}(\mathbf{k})$ is no longer homogenous but concentrates around the band inversion points, and their sign is not always positive (such as in 1CB) throughout the BZ, leading to relatively large fluctuation $\delta\mathcal{B}$ and deviation of $\mathbb{T}$ from the ideal value in LL.
@@
Figure 3: The distribution of Berry curvature $\mathcal{B}(\mathbf{k})$ and trace of Fubini-Study metric $\text{Tr}[g(\mathbf{k})]$ for the lowest two bands of continuum model and tight-binding model in the BZ. BZ is labelled by dashed hexagon. The fluctuation of Berry curvature $\delta\mathcal{B}$ are relatively larger in Chern bands compared LL, and thus $\mathbb{T}$ deviate more in Chern bands from the ideal LL.
@@
These differences between Chern bands of the tight-binding model and dispersive LL of the continuum model are inevitable, since we project out the higher energy degree of freedom in the continuum model and keep only $s$, $p_{-}$, $p_{+}$ Wannier states. Then the band inversion among them naturally lead to Berry curvature concentration. Especially for the 1CB, the band inversion between $p_{-}$ and $s$ contribute the negative $\mathcal{B}(\mathbf{k})$ around $\Gamma$ point and the inversion between $p_{-}$ and $p_{+}$ contribute the positive $\mathcal{B}(\mathbf{k})$ around $K$ and $K^{\prime}$ points. Remarkably, we will see that a non-Abelian state can occur even when the Berry curvature fluctuates strongly.
@@
Exact diagonalization.— To explore whether the non-Abelian state can appear in 1CB, we now study many-body physics at fractional filling in these Chern bands via numerical diagonalization. To make the many-body calculation tractable, we restrict our variational Hilbert space to that in which $N_{\text{uc}}$ electrons fill the LCB and $N_{e}-N_{\text{uc}}$ electrons remain in the 1CB where $N_{e}$ is the number of electrons. The electron-electron interaction Hamiltonian is obtained by projecting the realistic Coulomb interaction into the MLWFs and keeping the leading terms [53], which is defined as
@@
where $a$,$b$ = ($s$, $p_{-}$, $p_{+}$), $ij\right>$ means nearest-neighbor, $U$ and $V$ are the strength of onsite and nearest-neighbor interaction which is independent of orbital degree of freedom. We ignore the kinetic energy since these two lowest Chern bands are quite flat.
Fig. 4(a) display the many-body spectra at filling $\nu=1/3$ (namely, $1/3$ filling of LCB) as a function of crystal momentum $\mathbf{k}=k_{1}\mathbf{T}_{1}+k_{1}\mathbf{T}_{2}$, which is labeled as $k=k_{1}+N_{1}k_{2}$. Here $k_{i}=0,...,N_{i}-1$ ($i=1,2$) for system size $N_{\text{uc}}=N_{1}\times N_{2}$ with filled particle number $N_{e}=\nu N_{\text{uc}}$ and $\mathbf{T}_{i}$ are basis vectors of crystal momentum. The cluster size is chosen as $N_{\text{uc}}=4\times 6$, and $U=1$ and $V=0$. There are 3 nearly degenerate ground states well separated by a sizable energy gap from excited states. The approximate ground state degeneracy matches with the expected topological degeneracy of a fractional QH state on a torus. We further calculated different cluster sizes with all other parameters fixed and find the gap remains indicating its existence in the thermodynamic limit [53]. The lattice momenta of the degenerate ground states have linear indices $(0,8,16)$ and are in precise agreement with the generalized Pauli principle, which is the hallmark of FCI at $1/3$ filling [25]. To further confirm and distinguish FCI and other competing phases, we subsequently calculated the particle entanglement spectrum (PES) which encodes the information of the quasihole excitations [68, 69], by dividing the whole system into $N_{A}=3$ and $N_{e}-N_{A}$ particles. As shown in Fig. 4(b), we find that there is a clear entanglement gap separating the low-lying PES levels from higher ones for degenerate many body ground states. The number of PES levels below the gap exactly matches the typical counting of quasiparticle excitations resulting from the generalized Pauli principle of $1/3$ Laughlin state. These numerical results suggest that LCB resembles LLL.
@@
Fig. 4(c,d) show the many-body spectra at filling $\nu=3/2$ (that is, $1/2$ filling of 1CB assuming full spin polarization). The exact diagonalization is performed on finite-size torus with $U=1.3$ and $V=1$. For $N_{\text{uc}}=2\times 13$, the number of electrons occupying the 1CB is odd (13), while for $N_{\text{uc}}=2\times 14$, it is even (14). In these two cases, we observe two-fold and sixfold ground state quasi-degeneracies in Fig. 4(c) and 4(d), respectively. These are precisely the degeneracies expected for a Moore-Read state on the torus due to an even-odd effect [70, 71, 28], and the enhancement of gap indicates its existence in the thermodynamic limit. The lattice momenta at which these ground states occur also match the momenta of non-Abelian $\nu$ = 1/2 FCI based on the fractional quantum Hall-FCI folding scheme [25, 72]. We also calculate spectrum under the twist boundary condition to check the finite-size effect [53], where the momentum of single particle shifts to $k_{i}\rightarrow k_{i}+(\theta_{i}/2\pi)\mathbf{T}_{i}$ providing an effective way to scan all momenta in the BZ near continuously. We find the quasi-degenerate ground states remain well separated from the other low-energy excitation spectrum, indicating the robustness of the excitation gap [53].
@@
Discussions.— The method studied here can be further generalized to Chern bands with band inversion resembling higher LL. Concretely, an ideal local band inversion is introduced with a simple $\mathbf{k}\cdot\mathbf{p}$ model [75]

### v3

@@ Abstract
The pursuit of a lattice analogue for Landau levels has been a central theme in condensed matter physics. Although the correspondence between Chern bands and the lowest Landau level has been widely studied, a lattice realization of the first Landau level remains elusive. Here we construct a minimal lattice model that provides a concrete orbital description of both the lowest and first Landau levels. Using maximally localized Wannier functions with $s$, $p_{-}$, and $p_{+}$ orbital character, we develop a three-orbital model in which the two lowest Chern bands are flat and each carries a Chern number $\mathcal{C}=1$. The band topology arises from a sequence of ideal band inversions between Wannier states at the $\Gamma$ and $K$ points in momentum space, establishing an adiabatic connection between the atomic insulator limit and Landau level physics. Notably, many-body exact diagonalization reveals that the non-Abelian state can appear in the half-filled first Chern band. This construction can be further generalized to realize flat Chern bands analogous to higher Landau levels. Our results provide a new perspective on lattice analogues of Landau levels and may enable the exploration of fascinating topological phenomena at elevated temperatures.
@@
Introduction—The recent discovery of Abelian fractional Chern insulators (FCI) at zero magnetic field, dubbed the fractional quantum anomalous Hall state in moiré MoTe2 and pentalayer graphene [1, 2, 3, 4, 5] has generated intense interest in this new state of matter [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. The emergence of fractional topological states is attributed to the existence of flat Chern bands [20, 21, 22, 23, 24, 25, 26] with nearly ideal quantum geometry in moiré superlattice [27, 28, 29, 30, 31], which resembles the lowest Landau level (LLL). The exotic non-Abelian topological orders [32], such as Moore-Read states [33, 34, 35, 36] or Read-Rezayi state [37] are predicted to exist in the partially filled first Landau level (1LL). These states support non-Abelian quasiparticle excitations [34] and could be utilized as a platform for fault-tolerant quantum computation [38, 39]. This raises an interesting question [40, 41, 42, 43, 44, 45]: Is it possible to construct a lattice analogy of 1LL without explicitly invoking the magnetic field? This question is of significant importance because such a construction could pave the way for realizing non-Abelian fractionalization at elevated temperatures.
Here, rather than relying on the operator algebra to bridge Chern bands and LLs, we establish this connection via an adiabatic pathway from atomic limits to LL physics within the skyrmion Chern band model. We construct a projected three-band lattice model using maximally localized Wannier functions (MLWFs) with $s$, $p_{-}$, and $p_{+}$ orbital characteristics. The LL-like band topology in our model emerges through a sequence of ideal band inversions at high-symmetry points. We show using many-body exact diagonalization that the half-filled first Chern band (1CB) supports non-Abelian Moore-Read state. Our approach can be further extended to engineer flat Chern bands that mimic higher LLs. Notably, the corresponding lattice model is well-suited for implementation in cold atom systems, where the long-range Coulomb interaction is effectively projected onto a short-range density–density interaction.
@@
Intuition—As a starting point for constructing Chern bands in a lattice model, it is essential to understand the characteristics of the corresponding Wannier states. In earlier work, Qi constructed the one-dimensional MLWF in Chern bands that have a one-to-one mapping to the LLL wave functions in quantum Hall (QH) [23] of the form $\psi_{0,K_{y}}(x,y)\sim e^{iK_{y}y}H_{0}(x^{\prime})e^{-x^{\prime 2}/2l^{2}_{B}}$, where $x^{\prime}\equiv x-K_{y}\ell_{B}^{2}$. $H_{0}(x^{\prime})$ is the Hermite polynomial for $n=0$ LLL, and $H_{0}(x^{\prime})e^{-x^{\prime 2}/2l^{2}_{B}}=e^{-x^{\prime 2}/2l^{2}_{B}}$ is related to MLWF of Chern band and exhibits $s$-orbital–like character in one dimension. Intuitively, when this construction is extended to the 1LL, the MLWFs in the corresponding Chern band map to $H_{1}(x^{\prime})e^{-x^{\prime 2}/2l^{2}_{B}}=2x^{\prime}e^{-x^{\prime 2}/2l^{2}_{B}}$, which has a $p$-orbital–like structure. This suggests that, at a minimum, Wannier states with $s$- and $p$-orbital characteristics are necessary to construct a Chern band that captures the physics of the 1LL.
@@
Model and adiabatic connection—We begin with the skyrmion Chern band model in which itinerant electrons couple to a layer pseudospin skyrmion lattice. This model could well describe the consecutive topologically nontrivial flat bands in twisted homobilayer MoTe2 around twist angle $2^{\circ}$ [41, 46, 45, 47, 48, 49, 50, 51],
@@
Here, $\bm{\sigma}$ denotes the pseudospin operator associated with the layer degree of freedom, and $\mathbf{S}(\mathbf{r})=\mathbf{S}(\mathbf{r}+\mathbf{a}_{1,2})$ represents the periodic moiré potential and is naturally coupled to the pseudospin of the layer with strength $\mathcal{J}$. In the small twist-angle limit, the moiré potential becomes dominant, such that $\mathcal{J}\gg\hbar^{2}/ma^{2}$, and enforces the local alignment of the electron layer polarization with $\mathbf{S}(\mathbf{r})$. This alignment induces a pseudospin Berry phase. The effect can be made explicit through a position-dependent $SU(2)$ unitary transformation $\mathbf{U}(\mathbf{r})$, which rotates the pseudospin texture $\mathbf{S}(\mathbf{r})$ into $S(\mathbf{r})\hat{z}$, thereby introducing a gauge field $\mathbf{A}_{j}=(i\hbar/e)\mathbf{U}^{\dagger}\partial_{j}\mathbf{U}$. In the large-$\mathcal{J}$ limit, the system can be projected onto the low-energy manifold of locally layer-polarized electrons, yielding an effective Hamiltonian for the low-energy band as
@@
From such observations, we see that $\mathcal{H}_{\text{eff}}$ actually links effective atomic orbital and LL physics in the low-energy manifold. This is made clearly by calculating the band structure of $\mathcal{H}_{0}$ when we take $\mathbf{S}(\mathbf{r})=\mathbf{N}(\mathbf{r})/N^{\lambda}(\mathbf{r})$ with
@@
where $\hat{\mathbf{e}}_{j}=\left(i\alpha\sin\theta_{j},-i\alpha\cos\theta_{j},-1\right)/\sqrt{2}$, $\mathbf{q}_{j}=(4\pi/\sqrt{3}a)\left(\cos\theta_{j},\sin\theta_{j}\right)$, and the angles satisfy $\theta_{2}=\theta_{1}+2\pi/3$, $\theta_{3}=\theta_{1}+4\pi/3$, and $\theta_{j+3}=\theta_{j}+\pi$. This pseudospin texture can be viewed as a sum of three pseudospin spirals forming a triangular skyrmion lattice [55, 56, 57, 58], and the normalization is controlled by $\lambda$. When $\lambda=1$, $S(\mathbf{r})$ becomes spatially uniform, and the low energy bands of $\mathcal{H}_{0}$ are flat Chern bands with $\mathcal{C}=1$, as shown in Fig. 1(e). The lowest two Chern bands are dispersive LL, where their wave functions overlap with the corresponding flat LL exceed $99\%$ [59]. For simplicity, we label them as LLL and 1LL, respectively. Conversely, when $\lambda\rightarrow 0$, the minimum of $\mathcal{J}S(\mathbf{r})$ forms a moiré triangular lattice (see Fig. 1(g)), and three lowest bands in Fig. 1(a) are topologically trivial flat bands resembling the atomic insulator limit, which emerge from $s$, $_{-}, p+p_{+}$ orbital hopping on a triangular lattice.
@@
Figure 1: Band structures and adiabatic connection from atomic insulator to $L$. (a)-(e) The lowest few bands of $\mathcal{H}_{0}$ with different normalization parameter $\lambda$, where the overlap between Bloch state and Wannier function is labeled as different colors, blue for $s$ orbital, green for $p_{-}$ orbital, red for $p_{+}$ orbital and $\mathcal{C}$ denotes the Chern number of lowest three bands. The inset in (c) shows the Brillouin zone (BZ). (f) Band structure of the minimal three-orbital tight-binding model. (g) Real space distribution of $N(\mathbf{r})$. The parameters are $\mathcal{J}/(\hbar^{2}/ma^{2})=52\pi^{2}$, $\alpha=1$ and $N_{0}=0.28$.
@@
Fig. 1 display the evolution of the band structures from flat topologically trivial bands to flat Chern bands. As $\lambda$ changes from 0 to 1, the lowest three bands invert while their $s$, $_{-}, p+p_{+}$ orbital characteristics remain. To see the band inversion clearly, we graphically depict the overlap of the Bloch states with trial Wannier functions of $s$, $p_{-}$, and $p_{+}$ orbitals, symbolized by blue, green, and red, respectively. The lowest three bands first invert with the upper band at $M$ and their Chern numbers become $\mathcal{C}=[0,0,3]$ as shown in Fig. 1(b), then the second and third lowest bands invert at $K$ and switch Chern numbers to $\mathcal{C}=[0,2,1]$ as in Fig. 1(c), finally the lowest and second lowest bands invert at $\Gamma$ and Chern numbers becomes $\mathcal{C}=[1,1,1]$ as in Fig. 1(d,e). The lowest three Chern bands are isolated in Fig. 1(e). The orbital characteristics of the LLL and 1LL here are consistent with the mapping from the hybrid Wannier functions of Chern band to LL wave functions in QH [23].
@@
Wannier projection—Now we demonstrate that $s$, $_{-}, p+p_{+}$ orbitals constitute a complete subspace to describe LLL and 1LL. By constructing the MLWFs [63, 64, 65, 66] from the continuum model, a three-orbital model is developed to describe the lowest two Chern bands in Fig. 1(e), with the Wannier orbitals forming a triangular lattice. The minimal model includes hopping up to the fifteenth nearest neighbor, where its explicit form is in Supplemental Materials [59]. The band structure in Fig. 1(e) does not possess a local gap below which the total Chern number is zero. A set of frozen states is chosen to preserve the topology of the focused Chern bands and the band disentanglement [67] process is then performed to avoid Wannier obstruction [68]. We choose the lowest two Chern bands as frozen states, and get the minimal tight-binding model where only MLWFs with $s$, $_{-}, p+p_{+}$ orbital characteristics are included. The corresponding band structure is calculated in Fig. 1(f). The MLWFs of $s$, $_{-}, p+p_{+}$ orbitals are shown in Fig. 2. The lowest two bands projected from the LLL and 1LL are well reconstructed in Fig. 1(f), which are labeled LCB (lowest Chern band) and 1CB, respectively.
@@
Figure 2: The MLWFs of the lowest few bands in Fig. 1(e) with $s$, $_{-}, p+p_{+}$ orbital characteristics, and the white hexagon labels Wigner-Seitz cell.
@@
$(\delta\mathcal{B})^{2}$ $\equiv\frac{\Omega_{\text{BZ}}}{4\pi^{2}}\int_{\text{BZ}}d\mathbf{k}\left(\mathcal{B}(\mathbf{k})-\frac{2\pi\mathcal{C}}{\Omega_{\text{BZ}}}\right)^{2},$ (4)
$\mathbb{T}$ $\equiv\int_{\text{BZ}}d\mathbf{k}\left[\mathrm{Tr}(g(\mathbf{k}))\right],$ (5)
@@
where $\mathcal{B}(\mathbf{k})\equiv-2\mathrm{Im}(\eta^{xy})$ is the Berry curvature, $g(\mathbf{k})\equiv\mathrm{Re}(\eta^{\mu\nu})$ is the Fubini-Study metric, and $\eta^{\mu\nu}(\mathbf{k})\equiv\langle\partial^{\mu}u_{\mathbf{k}}|\left(1-|u_{\mathbf{k}}\rangle\langle u_{\mathbf{k}}|\right)|\partial^{\nu}u_{\mathbf{k}}\rangle$ is the quantum geometric tensor, $\mathcal{C}\equiv(1/2\pi)\int d^{2}\mathbf{k}\mathcal{B}(\mathbf{k})$, $\Omega_{\text{BZ}}$ is area of BZ. We plot the distribution of $\mathcal{B}(\mathbf{k})$ and $\text{Tr}[g(\mathbf{k})]$ of the BZ in Fig. 3. For LLL and 1LL, $\mathcal{B}(\mathbf{k})$ remain positive throughout the whole BZ and the distribution is quite homogeneous with relatively small fluctuation $\delta\mathcal{B}$, and $\mathbb{T}$ is almost ideal ($\mathbb{T}=2n+1$ for $n$th flat LL [73]). However, in the tight-binding model, $\mathcal{B}(\mathbf{k})$ is no longer homogenous but concentrates around the band inversion points, and their sign is not always positive (such as in 1CB) throughout the BZ, leading to relatively large fluctuation $\delta\mathcal{B}$ and deviation of $\mathbb{T}$ from the ideal value in LL.
@@
Figure 3: Distribution of Berry curvature $\mathcal{B}(\mathbf{k})$ and trace of Fubini–Study metric $\text{Tr}[g(\mathbf{k})]$ for the two lowest bands of both continuum model and tight-binding model across the Brillouin zone (BZ). The BZ is outlined by dashed hexagon. The bands derived from the continuum model are designated as LLL and 1LL, reflecting their highly ideal character and close resemblance to flat $L$s. In contrast, the bands from the tight-binding model are labeled LCB and 1CB. The Berry curvature fluctuation $\delta\mathcal{B}$ is noticeably larger in LCB and 1CB compared to their LLL and 1LL counterparts, resulting in a stronger deviation of the trace condition measure $\mathbb{T}$ from the ideal value.
@@
These differences between Chern bands of the tight-binding model and dispersive LL of the continuum model are inevitable, since we project out the higher energy degree of freedom in the continuum model and keep only $s$, $_{-}, p+p_{+}$ Wannier states. Then the band inversion among them naturally lead to Berry curvature concentration. Especially for the 1CB, the band inversion between $p_{-}$ and $s$ contributes the negative $\mathcal{B}(\mathbf{k})$ around $\Gamma$ point and the inversion between $p_{-}$ and $p_{+}$ contribute the positive $\mathcal{B}(\mathbf{k})$ around $K$ and $K^{\prime}$ points. Remarkably, we will see that a non-Abelian state can occur even when the Berry curvature fluctuates strongly.
@@
Exact diagonalization—To explore whether the non-Abelian state can appear in 1CB, we now study many-body physics at fractional filling in these Chern bands via numerical diagonalization. To make the many-body calculation tractable, we restrict our variational Hilbert space to that in which $N_{\text{uc}}$ electrons fill the LCB and $N_{e}-N_{\text{uc}}$ electrons remain in the 1CB where $N_{e}$ is the number of electrons. The electron-electron interaction Hamiltonian is obtained by projecting the realistic Coulomb interaction into the MLWFs and keeping the leading terms [59], which is defined as
@@
where $a$,$b$ = ($s$, $_{-}, p+p_{+}$), $\left<ij\right$> means nearest-neighbor, $U$ and $V$ are the strength of onsite and nearest-neighbor interaction. We adopt an isotropic approximation, neglecting minor variations in the interactions between different orbital pairs [59]. We also neglect the kinetic energy since these two lowest Chern bands are quite flat.
@@
Fig. 4(a) displays the many-body spectra at filling $\nu=1/3$ (that is, $1/3$ filling of LCB) as a function of crystal momentum $\mathbf{k}=k_{1}\mathbf{T}_{1}+k_{1}\mathbf{T}_{2}$, which is labeled $k=k_{1}+N_{1}k_{2}$. Here $k_{i}=0,...,N_{i}-1$ ($i=1,2$) for system size $N_{\text{uc}}=N_{1}\times N_{2}$ with filled particle number $N_{e}=\nu N_{\text{uc}}$ and $\mathbf{T}_{i}$ are basis vectors of crystal momentum. The cluster size is chosen as $N_{\text{uc}}=4\times 6$, and $U=1$ and $V=0$. There are 3 nearly degenerate ground states well separated by a sizable energy gap from excited states. The approximate ground state degeneracy matches with the expected topological degeneracy of a fractional QH state on a torus. We further calculated different cluster sizes with all other parameters fixed and found that the gap remains, indicating its existence in the thermodynamic limit [59]. The lattice momenta of the degenerate ground states have linear indices $(0,8,16)$ and are in precise agreement with the generalized Pauli principle, which is the hallmark of FCI at $1/3$ filling [25]. To further confirm and distinguish FCI and other competing phases, we subsequently calculated the particle entanglement spectrum (PES) which encodes the information of the quasihole excitations [74, 75], by dividing the whole system into $N_{A}=3$ and $N_{e}-N_{A}$ particles. As shown in Fig. 4(b), we find that there is a clear entanglement gap separating the low-lying PES levels from higher ones for degenerate many body ground states. The number of PES levels below the gap exactly matches the typical counting of quasiparticle excitations resulting from the generalized Pauli principle of $1/3$ Laughlin state. These numerical results suggest that LCB resembles LLL.
Fig. 4(c,d) show the many-body spectra at filling $\nu=3/2$ (that is, $1/2$ filling of 1CB assuming full spin polarization). We assume the energy separation between LCB and 1CB is much larger than the interaction ($U,V$), which prevents electrons in the filled LCB from gaining enough energy to be excited and participate in the many-body interactions involving the partially filled 1CB. Under these assumptions, the exact diagonalization is performed on finite-size torus with $U=1.3$ and $V=1$. For $N_{\text{uc}}=2\times 13$, the number of electrons occupying the 1CB is odd (13), while for $N_{\text{uc}}=2\times 14$, it is even (14). In these two cases, we observe two-fold and sixfold ground state quasi-degeneracies in Fig. 4(c) and 4(d), respectively. These are precisely the degeneracies expected for a Moore-Read state on the torus due to an even-odd effect [76, 77, 34], and the enhancement of gap indicates its existence in the thermodynamic limit. The lattice momenta at which these ground states occur also match the momenta of non-Abelian $\nu$ = 1/2 FCI based on the fractional QH-FCI folding scheme [25, 78]. We also find that under the twist boundary condition, the quasi-degenerate ground states remain well separated from the other low-energy excitation spectrum, indicating the robustness of the excitation gap [59].
@@
Discussions—The method studied here can be further generalized to Chern bands with band inversion resembling higher LL. Concretely, an ideal local band inversion is introduced with a simple $\mathbf{k}\cdot\mathbf{p}$ model [81]